# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v2

@@ 4.2 Exploring systematic patterns in judge models
In the previous section, we have seen that none of the judge models we considered were aligned as well with humans as the humans themselves. Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 10 points with the human-assigned scores. However, while this may limit – to some extent – the utility of using a judge models to get a perfect estimate of the exam-taker model’s capability on the benchmark, the judge models may still offer valuable insights to differentiate between different exam-taker models. If judges exhibit systematic biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
@@
To evaluate this, we compare the rankings assigned by each judge model to the nine exam-taker models by computing their Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking. The rankings are shown in Figure 3(a), with $\rho$ values in Appendix J. These results show that contains demonstrates the highest alignment with the human ranking, swapping the ranks of only two out of nine models. Notably, contains performs on par with JudgeLM 7B (Zhu et al., 2023), a language model fine-tuned specifically for evaluating language model responses. They are closely followed by GPT-4 Turbo and Llama-3 70B, the judges with the best alignment. Remarkably, while GPT-4 Turbo and Llama-3 70B rank the same three models in positions three, four, and five as the human judge, they do so in different orders. Several other judges have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones.
@@
Figure 3: (a) Contains and JudgeLM retain 67% of the human-assigned ranking, closely followed by GPT-4 Turbo and LLama3-70B. All judges struggle to distinguish between the poor-performing exam-taker models; (b) False positives and negatives across different judge models in descending order of human alignment. Both false negatives and false positives generally increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
@@ 5 Analysis
To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities.
5.1 Better aligned models: Recall gains, precision pains, and error spotlights
We first investigate the precision and recall of the judge models. We plot both – maintaining the ordering of Figure 1 – in Figure 4(a). It can be seen that the precision does not show any clear observable trend with increasing alignment, which can be further observed in Figure 3(b). The recall, on the other hand, shows an increasing trend, with more aligned models having comparatively fewer false negatives.
Figure 4: (a) Cohen’s kappa, precision, and recall for different judge models. Recall improves with increasing human alignment ($R^{2}$ = 0.98), though precision ($R^{2}$ = 0.0003) is not correlated with human alignment; (b) $\mathbf{\kappa}$ scores for each judge across different prompt templates. Except for GPT-4 Turbo, all judges struggle with too many detailed evaluation guidelines.
@@ 5.1 Better aligned models: Recall gains, precision pains, and error spotlights
Table 2: Error codes used to identify the types of errors made by exam-taker models when answering questions. The example question in this case is “Excluding Lady Jane Grey, who were the five monarchs of the House of Tudor?”, with the correct answer “Henry VII, Henry VIII, Edward VI, Mary I and Elizabeth I” (in any order).
@@ 5.2 Judge model sensitivity to prompt length and specificity
Next, we study the impact of the prompt on the ability of the judge models to perform an accurate assessment of exam-taker models, with two somewhat intertwined goals: 1) to study if the success of various judge models is related to the length of the prompt, and, 2) to study the degree to which the judgments of the judge models change with the specificity of the prompt.
We use four different prompt versions, varying in length and specificity. Each prompt instructs the judge models on how to evaluate responses, with complexity and detail increasing with the token count. The first two prompts (Without guidelines V1 and V2, with 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, with 245 and 301 tokens, respectively). All prompts are listed in Appendix K.
Figure 4(b) shows that Mistral 7B (Jiang et al., 2023), GPT-4 Turbo and Llama-3 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases. For this task, top performers’ (GPT-4 Turbo and Llama-3 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided. It can also be observed that only top performers appears to benefit from the more detailed instructions, with a slight upward trend, whereas the other models get less aligned with more instructions. This might be due to the less powerful judges not being able to follow many instructions in the prompt at the same time. Interestingly, in a follow-up experiment (see Appendix L), when the order of references provided to the judge models is shuffled, Figure 6(b) demonstrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models except Mistral 7B are more sensitive to the reference order given in the prompt.
@@ 5.3 Evaluating controlled responses
Figure 5: We observe that judge models remain robust when exam-taker models produce responses identical to the prompt. However, this robustness diminishes in the presence of hallucinated responses such as "Yes" and "Sure" from the exam-taker models. Additionally, when assessing the gold standard answer, judges do not consistently arrive at the correct judgement 100% of the time.
Next, we perform some simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses. For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatinm (the answer is thus always correct), while for the next three tests, the answer to be evaluated is always incorrect, with the dummy exam-taker model always responding with ‘‘Yes’’, and ‘‘Sure’’ for the second and the third tests, respectively, and simply repeating the question for the fourth test. Figure 5 shows that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B incorrectly evaluates a significant number of dummy answers, even though it shows a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)).
We hypothesise that when the answers are plausible but incorrect (e.g. if the question asks about the name of the author of a book, and the exam-taker model gives the name of the wrong author), most judges are able to identify them as being incorrect (by comparing it with the reference answer). However, the judges might get confused about what they are supposed to evaluate if the answer is completely unrelated to the question (such as the words ‘‘Yes’’ and ‘‘Sure’’). It is possible that in this situation a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
@@ 5.4 Leniency bias in judge models
Judge model $\kappa$ $P_{c}$ $P_{+}$
Gemma 2B 0.50 0.62 0.80
Llama-2 7B 0.66 0.68 0.36
Mistral 7B 0.72 0.75 0.75
Llama-2 13B 0.68 0.76 0.45
Llama-2 70B 0.80 0.79 0.94
Llama-3 8B 0.81 0.80 0.84
Llama-3 70B 0.84 0.84 0.79
GPT-4 Turbo 0.85 0.85 0.66
Figure 6: (a) Leniency bias across various judge models. To compute the leniency bias in judge models we estimate $P_{c}$ and $P_{+}$ for different judge models; (b) Consistency scores of each judge model for 3 random permutations of references in the prompt. The consistency score is the percentage of questions for which the judge model gives the same judgment for all 3 runs. We observe that the bigger models are more consistent as judges and less sensitive to reference orders in the prompt. More details are given in Appendix L.
To get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment. To do so, we assume that a judge assigns the correct judgment (i.e. same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias. We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results,444 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix M. and show them in Figure 6(a). We observe that $P_{+}$ for most models is significantly higher than $0.5$, indicating a tendency of the judge models to evaluate responses as ‘‘correct’’ when their evaluation criteria are not completely aligned with the provided instructions.
@@ 6 Conclusion
In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human evaluation and automated evaluation methods. By focusing on a relatively ‘clean’ evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm separately from the ambiguity and subjectivity in the task itself. We find that even in relatively straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs – also if they are specifically trained as judges, such as JudgeLM – and that even the best LLMs fail to meet the consistency of humans. Furthermore, though previous work commonly used percent agreement, we found that Cohen’s kappa ($\mathbf{\kappa}$) distinguishes judges much better. However, we observe that even judges with excellent $\mathbf{\kappa}$ scores may require further scrutiny. While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient approaches like JudgeLM and contains perform better when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and more systematic biases. In further analyses, we find that: 1) LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; 2) Judge models with lower alignment lack precision rather than recall; 3) GPT-4 Turbo and Llama-3 70B are generally robust across different prompts, but are difficult to ‘steer’ in their judgments; 4) Some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; 5) Judge models are better at detecting completely incorrect answers than partially incorrect ones.
@@
Overall, this work adds to the realm of LLM evaluation research by assessing judges within a clearly defined and objective framework. Our results highlight the utility of using some LLMs as judges but also urge caution in blindly trusting their judgments, even if they are found to be well-aligned with humans. For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Cohen’s kappa, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases. We further elaborate on the limitations of our work in Appendix A. In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability, and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
@@ Appendix A Limitations
In our work, we have shown potential pitfalls of using LLMs as judges, as well as the issues with commonly used metrics for quantifying their judgment quality. We have observed several trends and interesting phenomena regarding the behaviour of judge models in the LLM-as-a-judge paradigm, such as poor correlation or precision with alignment (§ 5.1), inability of smaller judge models to follow complex and nuanced instructions (§ 5.2), and tendencies of judge models to be lenient in their judgment when they are unsure (§ 5.4).
Broadly, our results show that caution and scrutiny are required in using LLMs as judges, but not all our specific conclusions about rankings may hold across different examples, different judge models and exam-taker model accuracies, and tasks. For example, on a small experiment with a different sample and different instructions, we observed that when we limited the questions to only those that have fewer than 10 or 20 references, there was a consistent 5-10 point drop in percent-alignment across all judge models, but the discriminative ability of the top-performing judges improved. In that example, in which the overall accuracy of the models was substantially lower than in the example in our main experiment, we also observed a closer match between Cohen’s kappa and percent alignment, suggesting that the accuracies of the exam-taker models being compared may play an important role in how well their scores match.
@@
Appendix B Model and dataset details

### v3

@@ 4.2 Exploring consistent patterns in judge models
In the previous section, we have seen that none of the judge models we considered were aligned with humans as well as the humans were aligned amongst themselves. Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 5 points with the human-assigned scores. However, while this may limit – to some extent – the utility of using a judge models to get a perfect estimate of the exam-taker model’s capability on the benchmark, the judge models may still offer valuable insights to differentiate between different exam-taker models. If judges exhibit consistent biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
@@
To evaluate this, we compare the rankings assigned by each judge model to the nine exam-taker models by computing their Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking. We show the rankings in Figure 3(a), with $\rho$ and corresponding $\sigma$ values in Appendix L. Most judge models have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones. Notably, the results show that several models that assign scores quite divergent from humans and have poor alignment on the sample level are very aligned in terms of the rankings they assign. Specifically, both contains and Mistral 7B, with Scott’s $\mathbf{\pi}$ values of 64 and 66, respectively, exhibit very high rank correlation with the human scores ($\rho$ 0.99 and 0.98, respectively, with $\sigma$ 0.02 and 0.03). With that, these judges perform on par with GPT-4 Turbo and outperform the better Llama judges – though with lower significance values – indicating that identifying which models are better should not be equated to assigning them the correct score.
@@
Figure 3: Judge rankings and true/false positives and negatives. (a) Assigned exam-taker model rankings assigned by highly human aligned judges. Contains stays closely to human-assigned rankings, as well as GPT-4 Turbo and Mistral 7B. (b) False positives and negatives across different judge models, in descending order of human alignment. Both false negatives and false positives increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
@@ 5 Analysis
To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities in the judges we investigate. Specifically, we study their precision and recall and error types (§ 5.1), their sensitivity to the instruction prompt prompt (§ 5.2), how they respond to controlled resposes of specific types (§ 5.3), and the extent to which they have a leniency bias (§ 5.4).
5.1 Better aligned models: Precision and recall gains with error spotlights
We first investigate the precision and recall of the judge models. Maintaining the ordering of Figure 1, we plot both in Figure 4(a). We can see that both precision and recall exhibit a moderate increasing trend as alignment increases. In Figure 3(b), we observe a similar pattern, though with a clearer picture on the distribution of false positives and negatives. Specifically, we see that the number of true positives is quite stable across many judges. The true negatives, instead, drop off quickly as the judge quality decreases, suggesting it is generally easier to judge answers that are correct.
Figure 4: Precision, recall and prompt sensitivity. (a) Recall and precision improve with increasing human alignment ($R^{2}$ = 0.31 and $R^{2}$ = 0.21, respectively). (b) Scott’s $\mathbf{\pi}$ scores for judges across different instructions.
@@ 5.1 Better aligned models: Precision and recall gains with error spotlights
Table 2: Error analysis for GPT-4 and Llama-3 70B judges. The example question is “Excluding Lady Jane Grey, who were the five monarchs of the House of Tudor?”, the correct answer “Henry VII, Henry VIII, Edward VI, Mary I and Elizabeth I” (in any order).
@@ 5.2 Judge model sensitivity to prompt length and specificity
Next, we study the impact of the prompt on the predictions of the judge models, to understand if the success of various judge models is related to the length of the prompt, and to study the degree to which the judgments of the judge models change with the specificity of the prompt. We use four different prompt versions, varying in length and specificity. The first two prompts (Without guidelines V1/V2, 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, 245 and 301 tokens, respectively). All prompts are listed in Appendix M.
Figure 4(b) shows that GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases. For this task, top performers’ (GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided. It can also be observed that only top performers appears to benefit from the more detailed instructions, with a slight upward trend, whereas the other models get less aligned with more instructions. This might be due to the less powerful judges not being able to follow many instructions in the prompt at the same time. In a follow-up experiment, we further investigate the impact of the order of the reference answers (for details, we refer to Appendix N). Figure 6(b) illustrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models – with the exception of Mistral 7B – are more sensitive to the reference order given in the prompt.
@@ 5.3 Evaluating controlled responses
Figure 5: Judge responses to dummy answers. We investigate how judge models respond to dummy answers. judge models remain robust when exam-taker models produce responses identical to the prompt (‘repeater’), but are less robust when the responses are "Yes" and "Sure". Even when the answer matches one of the reference answers verbatim (‘Gold answer’), judges do not always arrive at the correct judgement.
Next, we perform simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses. For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatim (the answer is thus always correct). For the next three tests, the answer is always incorrect. In the second and third tests, the dummy exam-taker model always responds with ‘‘Yes’’, and ‘‘Sure’’ for the second and third tests, respectively. For the fourth test, the evaluated answer is a repetition of the question.
In Figure 5, we can see that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B, incorrectly evaluate a significant number of dummy answers, even though they show a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)). We hypothesise that when the answers are plausible but incorrect (e.g. if the question asks about the name of the author of a book, and the exam-taker model gives the name of the wrong author), most judges are able to identify them as being incorrect (by comparing it with the reference answer). However, the judges might get confused about what they are supposed to evaluate if the answer is completely unrelated to the question (such as the words ‘‘Yes’’ and ‘‘Sure’’). It is possible that, in this situation, a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
@@ 5.4 Leniency bias in judge models
Judge model $\pi$ $P_{c}$ $P_{+}$
JudgeLM-7B 0.65 0.68 0.19
Mistral-7B 0.66 0.70 0.87
Llama2-70B 0.69 0.66 0.99
Llama2-13B 0.74 0.74 0.87
Llama3.1-8B 0.77 0.77 0.82
GPT-4 0.87 0.87 0.69
Llama3.1-70B 0.88 0.88 0.82
Llama3-70B 0.88 0.87 0.90
Figure 6: Leniency bias and answer consistency. (a) Leniency bias across various judge models. (b) Consistency score, defined as the percentage of questions for which the judge model gives the same judgment for three different answer orders.
Lastly, to get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment. To do so, we assume that a judge assigns the correct judgment (i.e. same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias. We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results555 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix O. and show them in Figure 6(a). We observe that $P_{+}$ for most models is significantly higher than $0.5$, indicating a tendency of the judge models to evaluate responses as ‘‘correct’’ when their evaluation criteria are not completely aligned with the provided instructions.
@@ 6 Conclusion
In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human judges as well as automated evaluation methods. By focusing on a clean evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm, separately from the ambiguity and subjectivity in the task itself. We find that even in straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs, such as Mistral 7B. GPT-4 Turbo, Llama-3.1 70B and Llama-3  70B, instead, are much better aligned, though they are still quite far from the alignment that humans have among each other. In some cases, despite their high alignment, their scores deviate from human scores with up to 5 points. Given the relative simplicity of the scenarios in which we deployed the judges, urging caution in using judge for more complex scenarios. Importantly, we noted that such patterns are virtually undetectable using the commonly deployed metric of percent aligned, which barely discrimates between the considered judges. We suggest that future work instead considers the more robust metric Scott’s $\mathbf{\pi}$, which allows to distinguish judges much better.
Next, we note that to discriminate between models, high alignment scores are not an absolute necessity. While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient and even the lexical matching method contains perform on par when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and score deviations. If the purpose of a study is to determine which model is better and not to estimate their actual scores, such approaches may thus be as suitable as the more expensive ones.
Lastly, we run a range of experiments to investigate judge models’ sensitivity to prompts, their precision and recall, their error types, how lenient they are, and how much they can be fooled by dummy answers. We find that LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; that judge models with lower alignment lack precision rather than recall, that better models are generally more robust across different prompts, but are difficult to ‘steer’ in their judgments; that some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; and that judge models are better at detecting completely incorrect answers than partially incorrect ones.
@@
Overall, this work adds to the realm of LLM evaluation research by assessing judges within a clearly defined and objective framework. Our results highlight the utility of using some LLMs as judges but also urge caution in blindly trusting their judgments, even if they are found to be well-aligned with humans. For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Scott’s $\mathbf{\pi}$, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases. We further elaborate on the limitations of our work in Appendix A. In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
@@ Appendix A Limitations
In our work, we have evaluated how 11 different LLMs fare as judges in a scenario in which judgements should be relatively straight-forward, and human alignment is high. As any study, our work has several limitations as well as directions that we did not explore but would have been interesting too. In this section, we discuss both.
Simplicity of the task
As mentioned in the introduction of our work, the scenario in which judges are used are typically much more complicated than the scenario that we focussed on. Specifically, judges are most often deployed in preference rankings (where two model responses are compared) or to judge complex answers that are difficult to automatically parse. In such tasks, human agreement is often low, making it challenging to judge the judges themselves. In our work, we have deliberately chosen for a simple task, in which human alignment is high. The main premise is, that if a judge does not perform well in this simple setup, caution is suggested also in more complex setups – if someone cannot do multiplication, why would they be able to solve ordinary differential equations. Given the poor understanding of which abilities of LLMs generalise in what dimensions, however, more studies are needed to understand how our results generalise to various other scenarios.
Human alignment
In an earlier version of this paper, due to the high cost of human annotations, we opted to select a single model for human annotation as we iteratively modified the exam taker prompt, few-shot examples, and guidelines. We selected the Llama2 7B for this purpose with a random sample of 600 questions. As this is only a single model, it is possible that our human alignment scores are biased because of that. After, we have therefore extended our results with another 600 human-annotated examples from Llama3.1 70B.
For Llama2 7B The average alignment among human evaluators had a Scott’s $\mathbf{\pi}$ of $96.36\pm 1.46$,and the average percent agreement was $98.33\%\pm 0.76\%$. For Llama3.1 70B, we noted that the average alignment among human evaluators had Scott’s $\mathbf{\pi}$ of $95.78\pm 0.30$,and the average percent agreement was $98.72\%\pm 0.10\%$. Given the similarity of these two numbers, we believe that these 1200 samples provide an adequate estimate. In the paper, we take the average.
Size of the judged samples
As each of the nine exam-taker models requires human annotations for each sample, we restricted our analysis to 400 samples in total. This sample size also allowed us to conduct manual annotations and error analysis within 75 human hours/200 GPU hours (see Appendix H) and give reliable confidence intervals while also providing the flexibility to compare a range of models. We were not able to increase the size due to the cost, but a statistical analysis (details provided in Appendix I) illustrated that the variance because of this sample size was very low.
Selection of judges
With our selection of judges, we have stuck to autoregressive judges that can be used off-the-shelve, as well as one LLM specifically trained to judge. They are – at the moment of writing – the ones that are most commonly used as LLM-judges, and we have tried to be comprehensive across size and family. Nevertheless, we acknowledge that there are other judges that we could have considered as well. As including more judges in – compared to including more exam-taker models– relatively straightforward because it requires only computational power, no manual annotation, we hope that others may evaluate their newly proposed judges using our setup as well.
Future work
@@
Appendix B A brief explanation of the theoretical issues with Cohen’s kappa
Cohen’s Kappa Coefficient (Cohen, 1960) is a statistic to measure inter-rater agreement for categorical responses. Cohen’s Kappa coefficient measures this agreement by computing the observed (percent) agreement between raters ($p_{o}$) and comparing it with the hypothetical probability of chance agreement ($p_{e}$), which is taken as a baseline, as follows:
$\kappa\equiv\frac{p_{o}-p_{e}}{1-p_{e}}$ (1)
In this equation, the chance agremeent $p_{o}$ constitutes the hypothetical probability that observed agreement occurred by chance, given the observed distributions of the considered raters, under the assumption that the probabilities the raters assign to the observed labels are independent. Specifically, it is defined as:
$p_{e}=\sum_{k}\widehat{p_{k12}}=^{ind}\sum_{k}\widehat{p_{k1}}\widehat{p_{k2}}=\sum_{k}\frac{n_{k1}}{N}\frac{n_{k2}}{N}=\frac{1}{N^{2}}\sum_{k}n_{k1}n_{k2},$ (2)
where $\widehat{p_{k12}}$ is the estimated probability that rater 1 and rater 2 will classify the same item as $k$, rewritten to $\widehat{p_{k1}}\widehat{p_{k2}}$ under the assumption that $p_{k1}$ and $p_{k2}$ are independent. The crux of the issue with this method of computation, is that $\widehat{p_{k1}}$ and $\widehat{p_{k2}}$ are estimated independently from the data. As such, the chance agreement adjusts for the observed average differences between raters, which is in fact part of what we intend to measure.
To address this issue, Scott’s Pi (Scott, 1955) instead defines the chance baseline under the assumption that the raters have the same distribution, which is estimated considering the joint distribution of rater 1 and rater 2, rather than considering them separately. It defines $p_{e}$ as:
$p_{e}=\sum_{k}\widehat{p_{k}^{2}}=\sum_{k}\sum_{k}(\frac{n_{k1}+n_{k2}}{2N})^{2}$ (3)
As such, contrary to Cohen’s Kappa, it captures differences surpassing the chance agreement if rater 1 and rater 2 were in fact equivalent. In other words, we compare against a baseline in which raters would be equivalent, and we measure how much they deviate from that.
Note that if the empirical distributions of rater 1 and rater 2 are the same, so will the values of Scott’s Pi and Cohen’s Kappa be. This also implies that for larger observed (percent) alignment values, the values for Cohen’s Kappa and Scott’s Pi will be closer.
Appendix C Model and dataset details